Find the tangent to the parabola y2 = 6x − 3 perpendicular to the line x + 3y = 7


1
Expert's answer
2021-11-18T14:10:36-0500

Tangent line to the parabola y2=6x-3,

"2yy'=6\\
y'=\frac{3}{y}\\"

And slope of the straight line x+3y=7 is "\frac{-1}{3}" .

"y=\frac{-1}{3}x+\frac{7}{3}"

Since, tangent to the parabola is perpendicular to the straight line.

Therefore, product of slope of two curves is -1.

"m_s.m_p=-1\\
\frac{3}{y}.\frac{-1}{3}=-1\\
y=1"

Now, by putting y=1 in the parabola gives the tangent point.

"1^2=6x-3\\
x=\frac{2}{3}"

Thus, tangent point is "(\frac{2}{3},1)" .

And, the equation of the tangent line is the line passing "(\frac{2}{3},1)" and having slope 3.

i.e.,

"y=mx+b\\
1=3.\frac{2}{3}+b\\
b=-1"

Therefore, equation of tangent line is y=3x-1.